An amusement park charges admission plus a fee for for each ride.
Admission plus 2 rides costs $10.Admission plus five rides cost $16.
What's the charge for admission?For each ride?

Answer by PBMathandscience(9)

Let A be the cost of admission, and R be the cost of a ride.
equation 1: A + 2*R = 10
equation 2: A + 5*R = 16
Subtract equation 1 from equation 2 to get equation 3:
equation 3: (A-A) + (5*R - 2*R) = 16-10
3*R = 6
R = 2
Substitute the value for R into equation 1:
A + 2*(2) = 10
A = 6
Check by substituting A & R into equation 2
6 + 5*2 = 6 + 10 = 16
So the admission price is $6 and the rides are $2 each.